In a quadltrl ab=ad=10 and BC=CD=5. What is value of angle BAD?(Without trigonometry)

Answer by ed42ptt(4)

Let me start this with AB = AD = 10 and BC = CD = 5
Prove AB = AD > BC = CD
If AB = AD and BC = CD then ACD = ABC => A(CD) = (AB)C => A(C)D = A(BC) =>
A(5) = (10)C => (10)C = (5)A => C = 5/10A => C = 1/2A thus 2C = A
(2C)B = (2C)D = 2(CB) = 2(CD) => Since A = 2C and BAD = B(2C)D =>
BC(2D) => BAD =>(5)(2)D = 10D => BAD => B(10) => BAD thus 10D = 10B = BAD and
D = B => BA = AD = 10 then B = 1 and D = 1 thus BAD = 10 because (1)(10) = 10
End line
Shape of a kite, boomerang, or a crossed quadrilateral
Not sure whether I did it correctly so let me know if this is what it should have been or not.
ED